Sam is three times as old as his sister. After six years, he is twice as old as his sister. How old is sam and how old is his sister?
Found 2 solutions by math_helper, ikleyn:
Answer by math_helper(2461)


a = sister's age
b = Sam's age

b = 3a (1)
b+6 = 2*(a+6) (2)

Substitute 3a for b from (1), into (2):
(3a)+6 = 2(a+6)
3a+6 = 2a+12
a = 6 --> b = 18
--
Ans: Sam is 18 years old and his sister is 6 years old.
---
Check: in 6 years: Sam will be 24, his sister will be 12, 24 = 2*12 (ok)




Answer by ikleyn(52810)

.
To tutor @math_helper

An advise from me :


    If you opened your file with the tag "pre",


    then do not forget to close it with "/pre".